An aqueous solution is made with the salt obtained from combining the weak acid acetic acid, CH3CO2H, and the weak base methylamine, CH2NH2. Is the solution acidic, basic, or neutral? Ka=1.8×10−5 and Kb=4.4×10−4
Answer #1

This is the salt formed by weak acid and weak base.

CH3COOH + CH3NH2 --------------> CH3COO-  CH3NH3+

PKa = -log Ka PKb=-log Kb PKw = -log Kw

PKa = -log (1.8 x 10-5) PKb = -log ( 4.4 X 10-4) PKw = - log 10-14

PKa = 4.74 PKb = 3.35 PKw = 14

pH = 1/2 ( PKw + PKa - PKb )

pH = 1/2 ( 14 + 4.74 - 3.35 )

pH = 7.69

Hence the solution made by this salt is basic solution.